OpenStudy (anonymous):

What is the slope of a line that passes through the point (−1, 1) and is parallel to a line that passes through (4, 6) and (−1, −4)?

OpenStudy (mathstudent55):

Parallel lines have the same slope. Find the slope of the parallel line. The slope of the line you are asked for is the same.

OpenStudy (mathstudent55):

Do you know how to find the slope of a line given two points on the line?

OpenStudy (anonymous):

I have less knowledge on slope as before, so not really anymore.

OpenStudy (mathstudent55):

Ok. It's not hard. Subtract the values of the y-coordinates. Subtract the values of the x-coordinates. then divide the first number by the second number.

OpenStudy (anonymous):

as in: x - x ----- y - y ?

OpenStudy (anonymous):

Or the other way around.

OpenStudy (mathstudent55):

No. y in numerator and x in denominator. The points are (4, 6) and (-1, -4) The y-coordinates are 6 and -4. Subtracting them you get: 6 - (-4) = 6 + 4 = 10

OpenStudy (anonymous):

Then we do the x coordinates like that?

OpenStudy (mathstudent55):

The x coordinates are 4 and -1. Subtracting you get: 4 - (-1) = 4 + 1 = 5 slope = (difference in y) / (difference in x) = 10/5 = 2

OpenStudy (mathstudent55):

Yes

OpenStudy (mathstudent55):

This is how slope is usually written as an expression. It means what we did above. Subtract the y's. Subtract the x's. Divide the subtraction of the y's by the subtraction of the x's. \(m = slope = \dfrac{y_2 - y_1}{x_2 - x_1} \)

OpenStudy (anonymous):

So, you get the y coordinates and subtract then that number divided by the x coordinates subtracted equal the slope? Well you just answered my question, thank you very much. (:

OpenStudy (mathstudent55):

That is it. You got it.

OpenStudy (mathstudent55):

Also, remember that parallel lines have the same slope.

OpenStudy (mathstudent55):

You're welcome.

OpenStudy (anonymous):

Awesome. :p